OpenStudy (anonymous):

*MEDAL and FAN* Because of the rainy season, the depth in a pond increases 3% each week. Before the rainy season started, the pond was 10 feet deep. What is the function that best represents the depth of the pond each week and how deep is the pond after 8 weeks? Round your answer to the nearest foot. Hint: Use the formula, f(x) = P(1 + r)x. A. f(x) = 10(0.03)x, 36 feet B. f(x) = 10(1.03)x, 14 feet C. f(x) = 10(1.3)x, 37 feet D. f(x) = 10(1.03)x, 13 feet

OpenStudy (anonymous):

P is the initial value. What is the initial depth?

OpenStudy (anonymous):

Not sure.

OpenStudy (anonymous):

It's given in the second sentence of the problem.

OpenStudy (anonymous):

10?

OpenStudy (anonymous):

Right. Now r is the percent change from week to week, EXPRESSED AS A DECIMAL. What is r?

OpenStudy (anonymous):

0.03

OpenStudy (anonymous):

Exactly. Now put these values into\[f\left( x \right) = P \left( 1+r \right)^{x}\]What do you get?

OpenStudy (anonymous):

f(x) = P(1 + 0.03) \[^{x}\]

OpenStudy (anonymous):

1 + 0.03 = 1.03

OpenStudy (anonymous):

Good. But you can simplify (1+0.03), right?

OpenStudy (anonymous):

Good. Put it all together.

OpenStudy (anonymous):

f(x) = 10(1.03)x

OpenStudy (anonymous):

Excellent. Now you're down to two possible answers. To calculate the depth after 8 weeks, calculate f(8). What do you get?

OpenStudy (anonymous):

Uhhh do I replace x with 8?

OpenStudy (anonymous):

In other words, calculate\[f \left( 8 \right) = 10\left( 1.03 \right)^{8} = ?\]

OpenStudy (anonymous):

14

OpenStudy (anonymous):

Not what I get. Try again. Remember, calculate 1.03^8 first, then multiply by 10.

OpenStudy (anonymous):

12.66770081387616 D:

OpenStudy (anonymous):

Perfect. Round it to get your answer. Well done!

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

You're welcome.